OpenStudy (anonymous):

a 2.0 kg mass is released from rest at the top of a plane inclined at a 20 degree above horizontal. The coefficient of kinetic friction between the mass and the plane is .20. What will be the speed of the mass after sliding 4.0 m along the plane.

OpenStudy (anonymous):

First draw a diagram

OpenStudy (anonymous):



OpenStudy (anonymous):

ok i've gotten that far

OpenStudy (anonymous):

alright sum the forces in the y and x directions

OpenStudy (anonymous):

so normal force in x is 20 * 9.8 cos 20 and y is 20*9.8 sin 20? is that right? Or what should i do i'm really confused.

OpenStudy (anonymous):

\[\sum F_x=ma_x=F_x-f_k\] \[\sum F_y=ma_y=N-W\]

OpenStudy (anonymous):